Cost of Units Completed and in Process

The charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production.

Work in Process—Assembly Department
Bal., 8,000 units, 65% completed 33,920 To Finished Goods, 184,000 units ?
Direct materials, 188,000 units @ $1.9 357,200
Direct labor 527,300
Factory overhead 205,120
Bal. ? units, 75% completed ?

Cost per equivalent units of $1.90 for Direct Materials and $3.90 for Conversion Costs.

a. Based on the above data, determine the different costs listed below.

If required, round your interim calculations to two decimal places.

1. Cost of beginning work in process inventory completed this period. $
2. Cost of units transferred to finished goods during the period. $
3. Cost of ending work in process inventory. $
4. Cost per unit of the completed beginning work in process inventory, rounded to the nearest cent. $

b. Did the production costs change from the preceding period?

c. Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?
Answer #1

1. Cost of beginning work in process inventory completed this period.

Beginning work in Progress Cost +Beginning WIP Equivalent Units * Cost per equivalent

=33920+8000*(1-65%)*3.90=$44840

2. Cost of units transferred to finished goods during the period.

Cost of beginning work in process inventory completed + Cost of Units started and Completed

Units started and Completed =

Ending Units of WIP=8000+188000-184000

Units started and Completed=188000-12000=176000 units

Cost of units transferred to finished goods during the period

=$44840+176000*(1.90+3.90)=$1065640

3. Cost of ending work in process inventory.

12000*1.90+12000*.75*3.90=$57900

4. Cost per unit of the completed beginning work in process inventory, rounded to the nearest cent.

=$44840/8000=$5.61

b. Did the production costs change from the preceding period?

Current year Production Cost =(1.90+3.90)=$5.80

Yes the production cost has increased

Assuming that the direct materials cost per unit did not change from the preceding period, did the conversion costs per equivalent unit increase, decrease, or remain the same for the current period?

Answer is yes icnreased

Increased by $5.80-$5.61=.20 $ per unit